OpenStudy (anonymous):

The equation y=1777x+27,153 can be used to predict the number y of gun deaths in USA x years after 2000, that is x=0 corresponds to 2000, x=2 corresponds to 2002, x=4 corresponds to 2004.and so on. 1. The predicted deaths in 2004 wil be? 20,045 2. The predicted death in 2007 will be?14,714 3. Deaths will be 12,937 in what year?2008 Are my answers right?

OpenStudy (saifoo.khan):

First one is wrong. D:

OpenStudy (saifoo.khan):

Second is wrong too. D:

OpenStudy (anonymous):

Dang outcha

OpenStudy (saifoo.khan):

1) 1777(4) + 27153

OpenStudy (anonymous):

1. 34,261?

OpenStudy (saifoo.khan):

Perfect. Can you solve the second one?

OpenStudy (anonymous):

1777(7)+27153=39,592?

OpenStudy (saifoo.khan):

\[\checkmark\]

OpenStudy (saifoo.khan):

For last one, 12,937 = 1777(x)+27153 Solve for x

OpenStudy (anonymous):

1995?

OpenStudy (saifoo.khan):

im getting x = -8. So, 2000 - 8 = 1992

OpenStudy (anonymous):

How are you getting the -8?

OpenStudy (saifoo.khan):

12,937 = 1777(x)+27153 12937 - 27153 = 1777x -14216/1777 = x x = -8

OpenStudy (anonymous):

Ohhh thanks!

OpenStudy (saifoo.khan):

:D